# Uncovering Weaknesses in Neural Code Generation

### v1

@@ IV-A2 Answer-related weaknesses
Figure 6: Example Showing How Match-Based Metrics May Misjudge Correct Diverse Responses.
IV-B Generated code-related weaknesses
@@
IV. Missing pivotal semantics. Some pivotal information in the intention prompt has not appeared in the generated code. Therefore, the generated code can not implement the prompt intention.
@@
Figure 7: An Example Where the LLMs Overlooks Critical Semantics of the Prompt.
V. Wrong API usage. This weakness pertains to the incorrect use of an existing API (i.e., correct API names but incorrect parameters) or the use of non-existent APIs, reflecting a type of syntactic error. Therefore, instances where incorrect APIs are employed correctly but cannot solve the problem indicated in the prompt are not classified under this category.
The issue of API recommendation has garnered significant attention, with several notable studies contributing to the field [19, 53, 66]. However, as illustrated by the simple example in Fig. 8, CodeGen2.5’s model erroneously invokes a non-existent API, unique(). Such errors, which can be described as hallucinations, are typically easy to identify using a compiler geared towards programming languages.
Figure 8: An example of LLM Generating Code with Non-existent API Calls.
Specially, we observed that PLMs struggle to generate high-quality code when prompts specify the usage of third-party libraries, such as Pandas, NumPy, matplotlib, Django, SQL and so on. In the 1,264 problematic cases, there are 671 ones involving third-party library usage errors (about 53.09%). On the other hand, we can see that the introduction of more rigorous benchmarks, such as BigCodeBench [67], is poised to enhance the evaluation of LLMs in code generation tasks that involve a variety of function calls. This, in turn, is expected to propel advancements in LLMs’ ability to generate code with accurate API usage even towards complex instructions.
VI. Lack of domain knowledge. Due to a lack of common knowledge in certain areas (e.g., IP address ranges, URL format or the definition of radial ticks), the generated code may contain errors or incomplete part.
@@
Figure 9: An Example Demonstrating Potential Domain Knowledge Deficiency in LLM during Code Generation.
VIII. Gold plating. This particular shortcoming, prevalent among LLMs like CodeGen2.5, CodeGeeX2 and ChatGPT, has been highlighted by our research. We use the term “gold plating” metaphorically to describe instances where an LLM produces code that is more elaborate or intricate than necessary to meet the specified prompt. During code generation tasks, these models are prone to introducing superfluous features or enhancements—elements that were never solicited by the prompts.
This issue is well-recognized within the field, since numerous studies adopting truncation strategies for refining the generated output [9, 39]. Popular methods prioritize the initial segments of generated code, truncating at predetermined stop sequences like ‘\$\backslash nclass$’, ‘\$\backslash ndef$’, and ‘\$\backslash nprint$. Such approaches, rather than preventing the generation of unnecessary functions or statements, act as a filtering process to remove them post-generation [9]. While these truncation strategies are adept at removing a substantial portion of the unnecessary code, they are often too generalized to discern all vital components with precision. Their one-size-fits-all nature can lead to the inadvertent exclusion of nuanced, context-specific elements that are essential for the code’s functional integrity. Furthermore, such crude filtering may not account for the logical flow and dependencies within the code, potentially disrupting the cohesive structure necessary for the code to operate as intended.
In our study, we employed the truncation methodology delineated in Nijkamp et al. [39], which offers a more robust analytical framework than that of Chen et al. [9]. Our investigation uncovered 1327 instances across three benchmark datasets where the truncated source code demonstrated merely a weak correlation with their respective original prompts. For instance, in the example in Fig.10, both CodeGen2.5 and CodeGeeX2 were found to populate the list ‘myList’ without accurately performing the ‘get elements’ action as stipulated by the prompt. In contrast, GPT-4 predominantly adheres to the prompt’s specifications, albeit it additionally assigns the retrieved elements to a pre-defined variable ‘selected_elements’.
Figure 10: A Demonstrative Example of Feature Overgeneration by LLMs.
IX. Code duplication. This represents another shortcoming that we’ve identified in LLMs, specifically within the scope of our evaluation of CodeGen2.5 and CodeGeeX2. These models occasionally produce redundant code in response to certain prompts, where the repetition may or may not be pertinent to the given task.
@@
Figure 11: Example of One Model Producing Repeated Code at Output Limit.
V Weakness distribution over datasets
@@
V-A Distribution on the bigger models.
@@
V-A1 Individual model view
Table IV delineates the proportion of each weakness type for each model across the three benchmarks. Firstly, GPT-4 distinctly outperforms the other models when it comes to the count of problematic cases, exhibiting fewer issues in both HumanEval+ and DS-1000 than CodeGeeX2 and CodeGen2.5. Within this comparison, CodeGeeX2 shows superior performance relative to CodeGen2.5, which corroborates the results documented in Table II.
@@ V-A1 Individual model view
TABLE IV: Distribution of Weakness Types Across Each of the Three Large-Scale Models (%).
@@ I II III IV V VI VII VIII IX
V-A2 Overview distribution: union and intersection
Union distribution In the context of the union view, we conceptualize the three large models as a singular, overarching virtual model. For each prompt-oriented record (i.e., the set of prompt-code pairs towards the same prompt), we meticulously examine the presence of weakness types across all three models. Any weakness type identified in at least one model is recorded on the overview weakness table, ensuring that no potential vulnerability is overlooked.
@@ V-A2 Overview distribution: union and intersection
Table V shows the Union distribution results over the three benchmarks. For better illustration, we highlight the weakness types with ratio of above 50% with light gray, and those more than 10% and less than 50% with light blue.
TABLE V: Distribution of Weakness Types Across the Three Large-Scale Models on Three Benchmarks (Union).A light gray background highlights those weakness types that exceed a 50% ratio within a benchmark, while a light blue background denotes weakness types with ratios ranging from 10% to 50%.
@@
Intersection distribution For intersection distribution view, we focus on revealing shared weaknesses across the three examined LLMs. By identifying these common vulnerabilities, we highlight areas of critical weakness which demand immediate attention. This approach not only underscores the most pressing weakness concerns but also paves the way for a more targeted and efficient source code review process in practical applications.
@@
Table VI illustrates the distribution of nine distinct weakness types across three datasets. To enhance visual clarity, we employ a color-coding system: the most salient weakness types are shaded dark gray, the secondary ones in light blue, and the tertiary with a light green background.
TABLE VI: Distribution of Weakness Types Across the Three Large-Scale Models on Three Benchmarks (Intersection). A light gray background emphasizes the most prevalent weakness type, while a light blue background signifies weakness types that are the second most prominent. The light green shade corresponds to the third most dominant types. Upon reviewing the ratios, these three categories fall within the ranges of [20%, $\infty$), [10%, 20%], and [4%, 6%], respectively.
. Benchmark Benchmark-based weaknesses Model-related weaknesses I. Inaccurate desc. (I.i,I.ii, I.iii) II. Overly complex prompts III. Grammar errors IV. Single answer biases V. Missing pivotal semantics VI. Wrong API usage VII. Lack of domain knowledge VIII.Gold plating IX. Code duplication CoNaLa (339) 26.84% (6.49%, 16.52%, 6.19%) 0.88% 0.00% 12.09% 5.60% 0.00% 0.59% 5.60% 0% HumanEval+ (115) 0.00% 0.00% 0.00% 0.00% 11.30% 0.00% 0.00% 0.00% 0.00% DS-1000 (277) 4.33% (0.00%, 0.00%, 4.33%) 5.05% 0.00% 0.00% 24.91% 0.00% 0.00% 0.36% 0%
@@
Observations and discussion Looking at the results in Table V and Table VI, we have the following four observations.
@@
  • Type V: Missing pivotal semantics, is the most prevalent weakness among the three larger models, whether we consider the intersection or union perspective. Specifically, Table VI shows that at least one model fails to include some crucial semantic elements in 65.78% of all CoNaLa samples, 66.09% in HumanEval+, and 80.51% in DS-1000. Similarly, Table V reveals that each of the three models exhibits this type of shortcoming in precisely 11.30% of HumanEval+ samples and 24.91% of DS-1000 samples.
@@
    This issue is most pronounced in DS-1000, possibly due to the complexity of its prompts. The prompts in DS-1000 are not only the longest, as shown in Table I, but each of them also incorporate an “event flow”—the original intent, failed attempts with outcomes, and the current status. While this provides a richer context, it may introduce additional noise that complicates the models’ understanding of the prompts’ true intentions.
@@
  • Type I: Inaccurate Description stands out as a significant weakness in CoNaLa. Both referenced tables indicate that all three models produce erroneous codes for the same 26.84% of samples, which have been identified as inaccurately described.
@@
    Compared to the other two datasets, CoNaLa’s prompts are the briefest in terms of token count (refer to Table I), suggesting simpler tasks. However, this brevity may result in a lack of explanatory detail and context, impeding the models’ ability to interpret ambiguous or incomplete semantics in CoNaLa prompts.
@@
  • DS-1000 exhibits two distinct prompt-related weaknesses: Type I: Inaccurate prompt description and Type II: Overly complex prompts. At least one model produces problematic code for 6.72% of all prompts, while all three models struggle with the same 4.33% of prompts labeled as inaccurate and another 5.05% categorized as overly complex.
  • Type VIII: Gold Plating The issue of Type VIII: Gold plating warrants significant attention, particularly in CoNaLa and DS-1000, with the problem being more severe in CoNaLa.
    This tendency might originate from a common underlying cause: the relatively simple nature of problems presented in CoNaLa’s prompts. LLMs appear prone to over-elaboration when faced with uncertain responses. This is supported by the presence of 5.42% Type VIII weaknesses in DS-1000, which has shorter average token lengths compared to HumanEval+. While the absence of Type VIII weaknesses in HumanEval+ could be linked to more complex variables—such as superior quality of prompt descriptions (i.e., no Type I issues)—it is noteworthy that responses in HumanEval+ tend to be more verbose.
V-B Distribution over two smaller models (Intersection)
We have also analyzed the intersection of weakness distributions for the two smaller models, with results presented in Table VII. For each dataset, the most frequent weaknesses are highlighted with a light gray background, and the second most frequent are denoted with a light blue background.
A comparison with the findings in Table VI on the large models reveals an overall increase in the distribution of most types of weaknesses for the smaller models, with the exception of Type I in the DS-1000 dataset.
@@ V-B Distribution over two smaller models (Intersection)
TABLE VII: Distribution of Weakness Types Across the Two Smaller Models on Three Benchmarks (Intersection)
@@
VI Manually Curating Prompts of CoNaLa
As discussed in Section V-A2, within the CoNaLa dataset, each of the three larger models encounters difficulties generating correct source code for the same subset of prompts—specifically, 26.84%—exhibiting Type I weaknesses (as detailed in Table VI). Additionally, a higher incidence of this weakness, approximately 40%, is noticeable in the code produced by the two smaller models (as shown in Table VII).
@@ VI Manually Curating Prompts of CoNaLa
Figure 12: Example Showing the Result of Manual Curating of One Prompt.
@@
We subsequently re-evaluated the three larger models using the updated prompts. The smaller models were excluded from this round of testing because some of these 90 records had been utilized in their fine-tuning process. Table VIII presents the Exact Match (EM) and CodeBLEU scores for both the original and revised prompts, clearly demonstrating enhanced performance across all three models on the refined benchmark. For a clearer depiction, we calculated the gains in EM and CodeBLEU and highlighted them in red. Notably, GPT-4 shows the greatest sensitivity to the prompt modifications, with gains of 400% in EM and 29% in CodeBLEU. While the EM for the code generated by CodeGen2.5 and CodeGeeX2 on the polished prompts remains unchanged—suggesting that the proportion of codes exactly matching the reference answer has not been affected—the CodeBLEU score improvements are nevertheless substantial, with gains of 12% for CodeGen2.5 and 7% for CodeGeeX2, respectively.
TABLE VIII: Comparisons of the Model Performance on the Original and Revised prompts.
@@
VII Discussion
VII-A Limitations and Opportunities
Paradigm of Effective Prompting for Code-PLMs is unclear. It is widely acknowledged that well-crafted prompts are essential for generating high-quality code. Yet, defining what constitutes an effective prompt for code-centric code-PLMs remains elusive. Based on our observations (see Table IV), inaccurate prompt description ranks among the top three types of weaknesses across all benchmarks and for each of the three big PLMs studied. Despite the existence of some recent work, the exploration of prompt patterns has predominantly been from a human perspective [54, 37]. These studies provide intriguing insights but are still preliminary; systematic evaluations across more diverse benchmarks and code-PLMs are imperative prior to their widespread adoption.
Moreover, there is an increasing trend in research aimed at improving existing prompts, employing techniques such as prompt clarification [25] and prompt decomposition [23, 14]. Yet, a definitive, ‘silver bullet’ approach remains elusive. The ongoing quest to perfect prompt engineering continues to inspire and challenge both researchers and practitioners in the field.
Fine-grained representation of pivot semantics in prompts lacks attention. While numerous code-generation methodologies have been directed towards modeling source code from various angles [47, 6, 63], there is a paucity of research on accurately capturing the semantic nuances conveyed by prompts.
Our analysis, as detailed in Tables VI, VII, and IV, indicates that the omission of crucial semantic content in prompts (Type V) is a prevalent issue. This trend persists across all five models evaluated, spanning both smaller and larger PLMs, and is especially evident in the complex benchmarks of HumanEval+ and DS-1000.

### v2

@@ 4.1.2. Answer-related weaknesses
Figure 6. Example Showing How Match-Based Metrics May Misjudge Correct Diverse Responses.
4.2. Generated code-related weaknesses
@@
Type IV. Missing pivotal semantics. Some pivotal information in the intention prompt has not appeared in the generated code. Therefore, the generated code can not implement the prompt intention.
@@
Figure 7. An Example Where the LLMs Overlooks Critical Semantics of the Prompt.
Type V. Wrong API usage. This weakness pertains to the incorrect use of an existing API (i.e., correct API names but incorrect parameters) or the use of non-existent APIs, reflecting a type of syntactic error. Therefore, instances where incorrect APIs are employed correctly but cannot solve the problem indicated in the prompt are not classified under this category.
The issue of API recommendation has garnered significant attention, with several notable studies contributing to the field (He et al., 2021; Wei et al., 2022; Zhou et al., 2022). However, as illustrated by the simple example in Fig. 8, CodeGen2.5’s model erroneously invokes a non-existent API, unique(). Such errors, which can be described as hallucinations, are typically easy to identify using a compiler geared towards programming languages.
Figure 8. An example of LLM Generating Code with Non-existent API Calls.
Specially, we observed that PLMs struggle to generate high-quality code when prompts specify the usage of third-party libraries, such as Pandas, NumPy, matplotlib, Django, SQL and so on. In the 1,264 problematic cases, there are 671 ones involving third-party library usage errors (about 53.09%). On the other hand, we can see that the introduction of more rigorous benchmarks, such as BigCodeBench (Zhuo et al., 2024), is poised to enhance the evaluation of LLMs in code generation tasks that involve a variety of function calls. This, in turn, is expected to propel advancements in LLMs’ ability to generate code with accurate API usage even towards complex instructions.
Type VI. Lack of domain knowledge. Due to a lack of common knowledge in certain areas (e.g., IP address ranges, URL format or the definition of radial ticks), the generated code may contain errors or incomplete part.
@@
Figure 9. An Example Demonstrating Potential Domain Knowledge Deficiency in LLM during Code Generation.
Type VIII. Gold plating. This particular shortcoming, prevalent among LLMs like CodeGen2.5, CodeGeeX2 and ChatGPT, has been highlighted by our research. We use the term “gold plating” metaphorically to describe instances where an LLM produces code that is more elaborate or intricate than necessary to meet the specified prompt. During code generation tasks, these models are prone to introducing superfluous features or enhancements—elements that were never solicited by the prompts.
This issue is well-recognized within the field, since numerous studies adopting truncation strategies for refining the generated output (Chen et al., 2021; Nijkamp et al., 2023). Popular methods prioritize the initial segments of generated code, truncating at predetermined stop sequences like ‘\$\backslash nclass$’, ‘\$\backslash ndef$’, and ‘\$\backslash nprint$. Such approaches, rather than preventing the generation of unnecessary functions or statements, act as a filtering process to remove them post-generation (Chen et al., 2021). While these truncation strategies are adept at removing a substantial portion of the unnecessary code, they are often too generalized to discern all vital components with precision. Their one-size-fits-all nature can lead to the inadvertent exclusion of nuanced, context-specific elements that are essential for the code’s functional integrity. Furthermore, such crude filtering may not account for the logical flow and dependencies within the code, potentially disrupting the cohesive structure necessary for the code to operate as intended.
In our study, we employed the truncation methodology delineated in Nijkamp et al. (Nijkamp et al., 2023), which offers a more robust analytical framework than that of Chen et al. (Chen et al., 2021). We observed several instances where the generated code failed to incorporate the core operation specified in the prompt. For instance, in the example in Fig.10, both CodeGen2.5 and CodeGeeX2 were found to populate the list ‘myList’ without accurately performing the ‘get elements’ action as stipulated by the prompt. In contrast, GPT-4 predominantly adheres to the prompt’s specifications, albeit it additionally assigns the retrieved elements to a pre-defined variable ‘selected_elements’.
Figure 10. A Demonstrative Example of Feature Overgeneration by LLMs.
Type IX. Code duplication. This represents another shortcoming that we’ve identified in LLMs, specifically within the scope of our evaluation of CodeGen2.5 and CodeGeeX2. These models occasionally produce redundant code in response to certain prompts, where the repetition may or may not be pertinent to the given task.
@@
Figure 11. Example of One Model Producing Repeated Code at Output Limit.
5. Weakness distribution over datasets
@@
5.1. Distribution on the bigger models.
@@
5.1.1. Individual model view
Table 4 delineates the proportion of each weakness type for each model across the three benchmarks. Firstly, GPT-4 distinctly outperforms the other models when it comes to the count of problematic cases, exhibiting fewer issues in both HumanEval+ and DS-1000 than CodeGeeX2 and CodeGen2.5. Within this comparison, CodeGeeX2 shows superior performance relative to CodeGen2.5, which corroborates the results documented in Table 2.
@@ 5.1.1. Individual model view
Table 4. Distribution of Weakness Types Across Each of the Three Large-Scale Models (%).
@@ I II III IV V VI VII VIII IX
5.1.2. Overview distribution: union and intersection
Union distribution. In the context of the union view, we conceptualize the three large models as a singular, overarching virtual model. For each prompt-oriented record (i.e., the set of prompt-code pairs towards the same prompt), we meticulously examine the presence of weakness types across all three models. Any weakness type identified in at least one model is recorded on the overview weakness table, ensuring that no potential vulnerability is overlooked.
@@ 5.1.2. Overview distribution: union and intersection
Table 5 shows the Union distribution results over the three benchmarks. For better illustration, we highlight the weakness types with ratio of above 50% with light gray, and those more than 10% and less than 50% with light blue.
Table 5. Distribution of Weakness Types Across the Three Large-Scale Models on Three Benchmarks (Union).
@@
Intersection distribution. For intersection distribution view, we focus on revealing shared weaknesses across the three examined LLMs. By identifying these common vulnerabilities, we highlight areas of critical weakness which demand immediate attention. This approach not only underscores the most pressing weakness concerns but also paves the way for a more targeted and efficient source code review process in practical applications.
@@
Table 6 illustrates the distribution of nine distinct weakness types across three datasets. To enhance visual clarity, we employ a color-coding system: the most salient weakness types are shaded dark gray, the secondary ones in light blue, and the tertiary with a light green background. Upon reviewing the ratios, these three categories fall within the ranges of [20%, $\infty$), [10%, 20%), and [4%, 10%), respectively.
Table 6. Distribution of Weakness Types Across the Three Large-Scale Models on Three Benchmarks (Intersection).
Benchmark Benchmark-based weaknesses Model-related weaknesses
CoNaLa (339) 26.84% (6.49%, 16.52%, 6.19%) 0.88% 0.00% 12.09% 5.60% 0.00% 0.59% 5.60% 0%
HumanEval+ (115) 0.00% 0.00% 0.00% 0.00% 11.30% 0.00% 0.00% 0.00% 0.00%
DS-1000 (277) 4.33% (0.00%, 0.00%, 4.33%) 5.05% 0.00% 0.00% 24.91% 0.00% 0.00% 0.36% 0%
@@
Observations and discussion. Looking at the results in Table 5 and Table 6, we have the following four observations.
@@
  • Type V: Missing pivotal semantics, is the most prevalent weakness among the three larger models, whether we consider the intersection or union perspective. Specifically, Table 5 shows that at least one model fails to include some crucial semantic elements in 65.78% of all CoNaLa samples, 66.09% in HumanEval+, and 80.51% in DS-1000. Similarly, Table 6 reveals that each of the three models exhibits this type of shortcoming in precisely 11.30% of HumanEval+ samples and 24.91% of DS-1000 samples.
@@
    Towards the three benchmark, this issue is most pronounced in DS-1000, possibly due to the complexity of its prompts. The prompts in DS-1000 are not only the longest, as shown in Table 1, but each of them also incorporate an “event flow”—the original intent, failed attempts with outcomes, and the current status. While this provides a richer context, it may introduce additional noise that complicates the models’ understanding of the prompts’ true intentions.
@@
  • Type I: Inaccurate Description is another issue that can not be ignored. It stands out as a significant weakness in CoNaLa. Both referenced tables indicate that all three models produce erroneous codes for the same 26.84% of samples, which have been identified as inaccurately described.
@@
    Compared to the other two datasets, CoNaLa’s prompts are the briefest in terms of token count (refer to Table 1), suggesting simpler tasks. However, this brevity may result in a lack of explanatory detail and context, impeding the models’ ability to interpret ambiguous or incomplete semantics in CoNaLa prompts.
    In addition to CoNaLa, DS-1000 also exhibits this type of weakness. Problematic code is produced by at least one model for 6.72% of the prompts, all of which have been identified as inaccurately described. More notably, 4.33% of all samples are consistently problematic across all three models due to their inaccuracy.
  • Type VIII: Gold Plating also warrants significant attention, particularly in CoNaLa and DS-1000, with the problem being more severe in CoNaLa.
    This trend may stem from a common underlying factor: the relatively simplistic nature of the problems presented, in contrast to the more complex ones found in HumanEval+. LLMs appear prone to over-elaboration when faced with uncertain responses. This is supported by the presence of 5.42% Type VIII weaknesses in DS-1000, which has shorter average token lengths compared to HumanEval+. While the absence of Type VIII weaknesses in HumanEval+ could be linked to more complex variables—such as superior quality of prompt descriptions (i.e., no Type I issues)—it is noteworthy that responses in HumanEval+ tend to be more verbose.
@@
5.2. Distribution over two smaller models (Intersection)
We have also analyzed the intersection of weakness distributions for the two smaller models, with results presented in Table 7. For each dataset, the most frequent weaknesses are highlighted with a light gray background, and the second most frequent are denoted with a light blue background.
A comparison with the findings in Table 6 on the large models reveals an overall increase in the distribution of most types of weaknesses for the smaller models, with the exception of Type I in the DS-1000 dataset.
@@ 5.2. Distribution over two smaller models (Intersection)
Table 7. Distribution of Weakness Types Across the Two Smaller Models on Three Benchmarks (Intersection)
@@
6. Manually Curating Prompts of CoNaLa
As discussed in Section 5.1.2, within the CoNaLa dataset, each of the three larger models encounters difficulties generating correct source code for the same subset of prompts—specifically, 26.84%—exhibiting Type I weaknesses (as detailed in Table 6). Additionally, a higher incidence of this weakness, approximately 40%, is noticeable in the code produced by the two smaller models (as shown in Table 7).
@@ 6. Manually Curating Prompts of CoNaLa
Figure 12. Example Showing the Result of Manual Curating of One Prompt.
@@
We subsequently re-evaluated the three larger models using the updated prompts. The smaller models were excluded from this round of testing because some of these 90 records had been utilized in their fine-tuning process. Table 8 presents the Exact Match (EM) and CodeBLEU scores for both the original and revised prompts, clearly demonstrating enhanced performance across all three models on the refined benchmark. For a clearer depiction, we calculated the gains in EM and CodeBLEU and highlighted them in red. Notably, GPT-4 shows the greatest sensitivity to the prompt modifications, with gains of 400% in EM and 29% in CodeBLEU. While the EM for the code generated by CodeGen2.5 and CodeGeeX2 on the polished prompts remains unchanged—suggesting that the proportion of codes exactly matching the reference answer has not been affected—the CodeBLEU score improvements are nevertheless substantial, with gains of 12% for CodeGen2.5 and 7% for CodeGeeX2, respectively.
Table 8. Comparisons of the Model Performance on the Original and Revised prompts.
@@
7. Discussion
7.1. Limitations and Opportunities
  • Paradigm of Effective Prompting for Code-PLMs is unclear. It is widely acknowledged that well-crafted prompts are essential for generating high-quality code. Yet, defining what constitutes an effective prompt for code-centric code-PLMs remains elusive. Based on our observations (see Table 4), inaccurate prompt description ranks among the top three types of weaknesses across all benchmarks and for each of the three big PLMs studied. Despite the existence of some recent work, the exploration of prompt patterns has predominantly been from a human perspective (White et al., 2023; Mishra et al., 2022). These studies provide intriguing insights but are still preliminary; systematic evaluations across more diverse benchmarks and code-PLMs are imperative prior to their widespread adoption.
    Moreover, there is an increasing trend in research aimed at improving existing prompts, employing techniques such as prompt clarification (Li et al., 2023c) and prompt decomposition (Khot et al., 2023; Dua et al., 2022). Yet, a definitive, ‘silver bullet’ approach remains elusive. The ongoing quest to perfect prompt engineering continues to inspire and challenge both researchers and practitioners in the field.
  • Fine-grained representation of pivot semantics in prompts lacks attention. While numerous code-generation methodologies have been directed towards modeling source code from various angles (Sun et al., 2019; Brockschmidt et al., 2019; Zhong et al., 2020), there is a paucity of research on accurately capturing the semantic nuances conveyed by prompts.
    Our analysis, as detailed in Tables 6, 7, and 4, indicates that the omission of crucial semantic content in prompts (Type V) is a prevalent issue. This trend persists across all five models evaluated, spanning both smaller and larger PLMs, and is especially evident in the complex benchmarks of HumanEval+ and DS-1000.